# Recovery of Sparse Graph Signals

### v1

@@ IV-B5 Algorithm
The algorithm for implementing the Graph-BNB-GIC method is initialized by the set $\mathcal{L}^{(0)}$ containing only the root node, which is characterized by the sets $(\mathcal{S}_{0},\mathcal{S}_{1})=(\emptyset,\emptyset)$. At each iteration, branching occurs from the leaf node with the smallest depth, where if there are multiple leaves with the smallest depth, then the leaf with the largest upper bound is selected. That is, the leaf node is selected by
@@
The branching is performed by the branch step outlined in Subsection IV-B2. Hence, to generate $\mathcal{L}^{(t+1)}$, all elements in $\mathcal{L}^{(t)}$ are duplicated excluding the selected leaf node. The sets of the selected node, $(\mathcal{S}_{0},\mathcal{S}_{1})$, are then split into two child nodes: ($(\{\mathcal{S}_{0}\cup k\},\mathcal{S}_{1})\}$ and $(\mathcal{S}_{0},\{\mathcal{S}_{1}\cup k\})$, where $k={\lvert\mathcal{S}_{0}\rvert}+{\lvert\mathcal{S}_{1}\rvert}$. The algorithm terminates when $L^{(t+1)}=U^{(t+1)}$, where $U^{(t+1)}$ and $L^{(t+1)}$ denote the maximum upper bound and maximum lower bound across all leaf nodes (nodes in $\mathcal{L}^{(t+1)}$), respectively. In this case, the maximum upper bound and maximum lower bound belong to the same leaf, and the set $\mathcal{S}_{1}$, which is associated with this leaf, represents the estimated support set (i.e. the algorithm output). Additionally, if the upper bound of a leaf is smaller than the global lower bound, $L^{(t+1)}$, then it is pruned (removed from the set $\mathcal{L}^{(t+1)}$) since the cost function associated with any of its potential descendants is assumed lower than its upper bound and that the global lower bound represents a cost function of a feasible solution. The proposed Graph-BNB-GIC method is summarized in Algorithm 3.
@@ IV-C Computational Complexity
The main goal of this work is to develop graph-based low-complexity methods for support set recovery of sparse graph signals. Specifically, we aim to design low-complexity methods that approximate the solution of the GIC optimization in (18), since a direct solution requires an exhaustive search with computational complexity that grows exponentially with the size of the graph. In particular, the exhaustive search requires $\sum_{j=1}^{{s}}\binom{{\lvert{\mathcal{V}}\rvert}}{j}$ floating-point operations (flops), as can be seen in (18). The computational complexity of the GM-GIC method from Algorithm 2 is primarily determined by the computations of the local GIC on the largest subset, and thus requires $O\left(\sum_{j=1}^{{s}}\binom{\max_{q}|\mathcal{D}_{q}|}{j}\right)$ flops. Hence, its complexity is significantly lower than those of the GIC methods, where $\forall q:\mathcal{D}_{q}\subseteq\mathcal{D}$ and from (29) and Theorem 2, ${\lvert\mathcal{D}\rvert}\ll{\lvert{\mathcal{V}}\rvert}$. The Graph-BNB-GIC method from Algorithm 3 is anticipated to have an even lower computational burden compared with the GM-GIC method, due to its iterative nature. However, as shown in Remark 2, in its worst-case scenario it requires an order of $O\left(\sum_{j=1}^{{s}}\binom{|\mathcal{D}|}{j}\right)$ flops. Consequently, even in the worst case, the computational complexity of the Graph-BNB-GIC method is significantly lower than that of the GIC method.
@@ V GFOC
In this section, we present the GFOC method, which is a new method for correcting existing estimated support sets of the sparse graph signal, $\hat{{\Omega}}$. This correction is obtained by optionally replacing the support set elements with nodes from their one-hop neighborhood, ${\mathcal{N}[\hat{{\Omega}}]}$. The initial estimated support set, $\hat{{\Omega}}$, can be obtained by any sparse recovery method, including standard techniques (such as the OMP [31] and Lasso [23] methods), as well as the GM-GIC and Graph-BNB-GIC methods proposed in Subsection II-C.
@@ V-A Method
${\mathcal{N}(k)}\stackrel{{\scriptstyle\triangle}}{{=}}\{{\mathcal{N}_{\scriptstyle{1}}[k]}\setminus k\}=\{m\in{\mathcal{V}}:\delta(k,m)=1\},$ (51)
where ${\mathcal{N}_{\scriptstyle{1}}[k]}$ is defined in (5) (with $\Delta=1$). If the GIC increases, then the method propagates by replacing the $k$th support set element with its neighbor, i.e. the element in (51) that maximizes the GIC. Otherwise, the $k$th element stays in the estimated support. It is noted that if the GFOC method input is infeasible, i.e. ${\lvert\hat{{\Omega}}\rvert}>s$, then the input can be corrected by selecting the indices associated with the $s$ largest values in $\hat{{{\bf{x}}}}$ where $\hat{{{\bf{x}}}}={({{\bf{H}}}_{\hat{{\Omega}}}^{T}{{\bf{H}}}_{\hat{{\Omega}}}^{\leavevmode\nobreak\ })^{-1}{{\bf{H}}}_{\hat{{\Omega}}}^{T}}$. The GFOC method is summarized in Algorithm 4. In addition, the following simple example illustrates the process of the GFOC method for a $9$-node graph.
@@
2Define ${obj}({\Omega})={\|{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{{\Omega}}}}}{{\bf{y}}}\|}^{2}-{\rho({\lvert{\Omega}\rvert})}$
@@ V-B Discussion
[…]
${\mathcal{N}_{\scriptstyle{\Psi}}(k)}\stackrel{{\scriptstyle\triangle}}{{=}}\{{\mathcal{N}_{\scriptstyle{\Psi}}[k]}\setminus k\}=\{m\in{\mathcal{V}}:0<\delta(k,m)\leq\Psi\}.$ (52)
@@ VI Simulations
In this section, we evaluate the performance of the support set recovery methods developed in Section IV and Section V in numerical simulations. The simulation setup and compared methods are presented in Subsection VI-A. In Subsections VI-B and VI-C, the performance of the methods is evaluated based on their accuracy and computing time, respectively.
VI-A Setup
VI-A1 Model
We consider the support set recovery problem in (16) with the sparse dictionary matrix selected as the graph filter defined by (6)-(9). The graph shift operator used to construct the graph filter is the graph Laplacian matrix [3], i.e. ${\bf{S}}={\bf{L}}$. For simplicity, the graph Laplacian matrix is normalized such that its largest eigenvalue is $12$. Additionally, the measurement vector ${{\bf{y}}}$ used for (16) is defined in (11). In our simulations we use the two-dimensional, square grid graph and the Erdős-Rényi graph [70], where for each graph, the edge weights are drawn from a uniform distribution $U([0,1])$. The graph filter coefficients are set to $1$, $h_{j}=1$, $j=0,\ldots{\Psi}$. The values assigned to the non-zero elements in the sparse graph signal, ${{\bf{x}}}$, are drawn from a standard normal distribution. The filtered signal, ${{\bf{H}}}{\bf{x}}$, is normalized to a specified signal-to-noise ratio (SNR) according to $\text{SNR}={\|{{\bf{H}}}{{\bf{x}}}\|}^{2}/({\lvert{\mathcal{V}}\rvert}\sigma_{n}^{2})$, where the noise covariance matrix satisfies ${\bf{R}}=\sigma_{n}^{2}{\bf{I}}$. Unless otherwise specified, the SNR and noise variance are set to $\text{SNR}=50$ and $\sigma_{n}=0.01$.
VI-A2 Methods
@@
  • The Lasso method [23];
@@
  • The GM-GIC method in Algorithm 2;
  • The Graph-BNB-GIC method in Algorithm 3 (denoted as G-BNB);
@@
  • The GFOC support set correction method in Algorithm 4, with the estimated support set input from each of the above mentioned methods; The corrected outputs are denoted as 1) OMP¯$^{c}$, 2) $\overline{\text{Lasso}}^{c}$, 3) $\overline{\ell_{1}-BnB}^{c}$, 4) GM-GIC¯$^{c}$, 5) $\overline{\text{G-BNB}}^{c}$, 6) GIC¯$^{c}$.
In practice, the Lasso and $\ell_{1}$-BNB methods directly solve the sparse recovery problem in (14). As shown in Section II, this is equivalent to solving (16). The tuning factors for the different methods are stated as follows. The Lasso and $\ell_{1}$-BNB regularization parameter is set to $0.01$. Additionally, the GIC penalty function, ${\rho({\lvert{\Omega}\rvert})}$, used for the GM-GIC, GIC, G-BNB, GIC, and GFOC methods, is chosen to be the AIC penalty function [66], i.e. we set $p({\lvert{\Omega}\rvert})=2{\lvert{\Omega}\rvert}$. For the GM-GIC method, the pre-screening threshold $\zeta$ in (35) is set to $\sigma_{n}$. For each presented scenario, the performancehas been evaluated by at least $10^{3}$ Monte Carlo simulations.
VI-B support set recovery Accuracy
In this subsection, we study the accuracy of the support set recovery by the methods in Subsection VI-A2, where the accuracy is measured by the F-score classification metric [71]. The F-score compares between the the true support set, ${\Omega}_{t}$, and the estimated support, $\hat{{\Omega}}$, which is obtained by the different methods in Subsection VI-A2. The F-score is given by
@@
where $t_{p}\stackrel{{\scriptstyle\triangle}}{{=}}{\lvert{\Omega}_{t}\cap\hat{{\Omega}}\rvert}$ is the number of true-positives, $f_{p}\stackrel{{\scriptstyle\triangle}}{{=}}{\lvert\{{\mathcal{V}}\setminus{\Omega}_{t}\}\cap\hat{{\Omega}}\rvert}$ is the number of false-positives, and $f_{n}\stackrel{{\scriptstyle\triangle}}{{=}}{\lvert{\Omega}_{t}\cap\{{\mathcal{V}}\setminus\hat{{\Omega}}\}\rvert}$ is the number of false-negatives. The F-score is between 00 and $1$, where $1$ indicates perfect support set recovery, i.e. $\hat{{\Omega}}={\Omega}$.
VI-B1 Case 1 - Square Grid Graph with $100$ Nodes
We consider the two scenarios presented in Fig. 4:
@@ VI-B1 Case 1 - Square Grid Graph with $100$ Nodes
    Scenario 1 (Dispersed Support): In this scenario, the signal support set is located on the edges of the square grid graph. For the considered $100$-node square grid graph, the node positions are ${\Omega}_{t}=\{1,10,91,100\}$ (shown in red in Fig. 4). This scenario describes a situation where the locations of the non-zero elements of the sparse signal are far apart across the graph.
@@
    Scenario 2 (Localized Support): In this scenario, the support set of the signal is concentrated in neighboring nodes (here, in the center of the square grid graph). For the considering $100$-node square grid graph, the node positions are ${\Omega}_{t}=\{45,46,55,56\}$) (shown in blue in Fig. 4). This scenario describes a situation where the non-zero elements are clustered in close proximity within the graph structure.
Figure 4: Square grid graph with ${\lvert{\mathcal{V}}\rvert}=100$. The nodes of Scenario 1, ${\Omega}_{t}=\{1,10,91,100\}$, and Scenario 2, ${\Omega}_{t}=\{45,46,55,56\}$, are colored in red and blue, respectively.
(a) Scenario 1
(b) Scenario 2
Figure 5: Square grid graph: the F-score of the different methods versus the graph filter degree, ${\Psi}$, for the two scenarios from Fig. 4.
The F-score results for Scenario 1 and Scenario 2 are presented in Figs. 5.a and 5.b, respectively, versus the graph filter degree. Figure 5.a shows that for Scenario 1, the F-score for the GM-GIC, G-BNB, and OMP methods is similar for any graph filter degree and achieves the benchmark score of the GIC method. This can be explained by the fact that a dispersed support set is easy to recover. Additionally, the F-score decreases moderately as the graph filter degree increases (from $0.95$ to $0.75$). In contrast, the F-score of the Lasso and the $\ell_{1}$-BNB methods significantly decreases as the graph filter degree increases (e.g. the F-score of the $\ell_{1}$-BNB method ranges from $0.8$ to $0.18$). For this case, the proposed GFOC method significantly improves the $\ell_{1}$-BNB method’s F-score for any graph filter degree, and improves the F-score of the Lasso method for $\Psi=1,2,3$. Figure 5.b shows that for Scenario 2 of the localized support, only the GM-GIC achieves F-scores comparable to the optimal GIC method. The G-BNB method is the closest behind, with F-scores lower by at most $0.15$. The corrected version $\overline{\text{G-BNB}}^{c}$ method performs even better, with F-scores lower than those of the GM-GIC method by at most $0.1$. The Lasso method obtains equivalent results to the GM-GIC method for $\Psi=1,2$, but provides poor results for graph filter degrees larger than $2$. Both the OMP and $\ell_{1}$-BNB methods exhibit a significant decrease in F-score as the graph filter degree increases. Finally, the improvement obtained by the proposed GFOC method is the most significant when applied to the OMP technique.
Figure 6 presents the F-score versus the SNR for Scenario 2, where the graph filter degree is set to ${\Psi}=4$. In this scenario, the GIC, GM-GIC, and G-BNB  methods have the highest F-scores (ranging from approximately $0.4$ to $0.85$), which increases as the SNR increases. The corrected version of the OMP method, OMP¯$^{c}$, improves the OMP F-score by about $0.15$, compared to the original OMP. Additionally, the corrected versions of the $\ell_{1}$-BNB, and $\overline{\ell_{1}-BnB}^{c}$ methods show similar results to those from the OMP¯$^{c}$ method for high SNRs. The Lasso method and its corrected version show poor results across all observed SNRs.
Figure 6: Square grid graph: the F-scores of the different methods versus the SNR with ${\Omega}_{t}$ from Scenario $2$, and ${\Psi}=4$.
VI-B2 Case 2- Erdős Réyni Graph with $70$ Nodes
We consider the graph in Fig. 7, which is a realization of an Erdős-Rényi graph with the probability of $0.12$ that two nodes are connected. The signal support set has been selected uniformly from the following two cases:
  1. -
    Scenario 3 (Mixed Dispersed-Localized Support): In this scenario, two groups of two nodes are selected as follows. First, two nodes are selected at random from the Erdős Réyni graph in Fig. 7. Then, for each node, one of its neighbors is randomly chosen and added to the support set. This process is repeated until four different nodes are selected. This scenario describes a situation where the non-zero elements are clustered in two distanced groups, where the nodes in each group are in close proximity.
  2. -

### v2

@@ IV-B5 Algorithm
The algorithm for implementing the gbnb method is initialized by the set $\mathcal{L}^{(0)}$ containing only the root node, which is characterized by the sets $(\mathcal{S}_{0},\mathcal{S}_{1})=(\emptyset,\emptyset)$. At each iteration, branching occurs from the leaf node with the smallest depth, where if there are multiple leaves with the smallest depth, then the leaf with the largest upper bound is selected. That is, the leaf node is selected by
@@
The branching is performed by the branch step outlined in Subsection IV-B2. Hence, to generate $\mathcal{L}^{(t+1)}$, all elements in $\mathcal{L}^{(t)}$ are duplicated excluding the selected leaf node. The sets of the selected node, $(\mathcal{S}_{0},\mathcal{S}_{1})$, are then split into two child nodes: ($(\{\mathcal{S}_{0}\cup k\},\mathcal{S}_{1})\}$ and $(\mathcal{S}_{0},\{\mathcal{S}_{1}\cup k\})$, where $k={\lvert\mathcal{S}_{0}\rvert}+{\lvert\mathcal{S}_{1}\rvert}$. The algorithm terminates when $L^{(t+1)}=U^{(t+1)}$, where $U^{(t+1)}$ and $L^{(t+1)}$ denote the maximum upper bound and maximum lower bound across all leaf nodes (nodes in $\mathcal{L}^{(t+1)}$), respectively. In this case, the maximum upper bound and maximum lower bound belong to the same leaf, and the set $\mathcal{S}_{1}$, which is associated with this leaf, represents the estimated support set (i.e. the algorithm output). Additionally, if the upper bound of a leaf is smaller than the global lower bound, $L^{(t+1)}$, then it is pruned (removed from the set $\mathcal{L}^{(t+1)}$) since the cost function associated with any of its potential descendants is assumed lower than its upper bound and that the global lower bound represents a cost function of a feasible solution. The proposed gbnb method is summarized in Algorithm 3.
@@ IV-C Computational Complexity
The main goal of this work is to develop graph-based low-complexity methods for support set recovery of sparse graph signals. Specifically, we aim to design low-complexity methods that approximate the solution of the gic optimization in (18), since a direct solution requires an exhaustive search with computational complexity that grows exponentially with the size of the graph. In particular, the exhaustive search requires $\sum_{j=1}^{{s}}\binom{{\lvert{\mathcal{V}}\rvert}}{j}$ floating-point operations (flops), as can be seen in (18). The computational complexity of the gmgic method from Algorithm 2 is primarily determined by the computations of the local gic on the largest subset, and thus requires $O\left(\sum_{j=1}^{{s}}\binom{\max_{q}|\mathcal{D}_{q}|}{j}\right)$ flops. In particular, when the set $\mathcal{D}$ cannot be partitioned, the complexity becomes $O\left(\sum_{j=1}^{{s}}\binom{\mathcal{D}}{j}\right)$. Hence, its complexity is significantly lower than those of the gic methods, where $\forall q:\mathcal{D}_{q}\subseteq\mathcal{D}$ and from (31) and Theorem 4, ${\lvert\mathcal{D}\rvert}\ll{\lvert{\mathcal{V}}\rvert}$. The gbnb method from Algorithm 3 is anticipated to have an even lower computational burden compared with the gmgic method, due to its iterative nature. However, as shown in Remark 2, in its worst-case scenario it requires an order of $O\left(\sum_{j=1}^{{s}}\binom{|\mathcal{D}|}{j}\right)$ flops. Consequently, even in the worst case, the computational complexity of the gbnb method is significantly lower than that of the gic method. This analysis highlights that the main advantage of the proposed methods lies in their reliance on the size of the set $\mathcal{D}$ for complexity considerations, rather than the size of the entire node set, $\mathcal{D}$, thereby enabling efficient computation even in large-scale applications.
@@
In this section, we present the gfoc method, which is a new method for correcting existing estimated support sets of the sparse graph signal, $\hat{{\Omega}}$. This correction is obtained by optionally replacing the support set elements with nodes from their one-hop neighborhood, ${\mathcal{N}[\hat{{\Omega}}]}$. The initial estimated support set, $\hat{{\Omega}}$, can be obtained by any sparse recovery method, including standard techniques (such as the omp [31] and Lasso [23] methods), as well as the gmgic and gbnb methods proposed in Section IV.
@@ V-A Method
${\mathcal{N}(k)}\mathrel{{}\mathop{=}\limits^{\triangle}}\{{\mathcal{N}_{\scriptstyle{1}}[k]}\setminus k\}=\{m\in{\mathcal{V}}:\delta(k,m)=1\},$ (53)
where ${\mathcal{N}_{\scriptstyle{1}}[k]}$ is defined in (5) (with $\Delta=1$). If the gic increases, then the method propagates by replacing the $k$th support set element with its neighbor, i.e. the element in (53) that maximizes the gic. Otherwise, the $k$th element stays in the estimated support. It is noted that if the gfoc method input is infeasible, i.e. ${\lvert\hat{{\Omega}}\rvert}>s$, then the input can be corrected by selecting the indices associated with the $s$ largest values in $\hat{{{\bf{x}}}}$ where $\hat{{{\bf{x}}}}={({{\bf{H}}}_{\hat{{\Omega}}}^{T}{{\bf{H}}}_{\hat{{\Omega}}}^{\leavevmode\nobreak\ })^{-1}{{\bf{H}}}_{\hat{{\Omega}}}^{T}}$. The gfoc method is summarized in Algorithm 4. In addition, the following example illustrates the process of the gfoc method for a $9$-node graph.
@@
2Define ${obj}({\Omega})={\|{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{{\Omega}}}}}{{\bf{y}}}\|}^{2}-{\rho{\lvert{\Omega}\rvert}}$
@@ V-B Discussion
The gfoc method is derived based on the gic in (18) and the analysis in Section III, derived following Assumptions A.1-A.2. Specifically, based on the discussion around (31), it can be inferred that when estimating the support set, if some errors occur, they are likely to be caused by the accidental inclusion of nodes in close proximity to the elements in the true signal support set ${\Omega}_{t}$. These mistakenly included nodes may be in addition to, or instead of, some of the elements in ${\Omega}_{t}$. In response to this possible error, Algorithm 4 proposes examining the immediate neighborhood of elements within the estimated support set to potentially enhance solutions wrt the gic. It is noted that using the gfoc method may only improve the performance compared to the method’s input wrt to the gic. Moreover, it can be verified from Algorithm 4 that the number of GIC computations required is limited by ${s}d_{max}$, where $d_{max}$ represents the maximum degree of the graph. Now, due to Assumption A.1, every node in the graph is connected to a small number of neighbors. Hence, we can assume that $d_{\max}$ is relatively low, i.e. $d_{\max}\ll{\lvert{\mathcal{V}}\rvert}$. Thus, the computational complexity of this approach is expected to be low. It should be noted that the support of the output graph signal obtained by the gfoc method can be smaller than the input support set, as a single neighbor may replace two elements. The gfoc method can be easily extended to the graph $\Psi$ order correction by replacing (53) with
$\{{\mathcal{N}_{\scriptstyle{\Psi}}[k]}\setminus k\}=\{m\in{\mathcal{V}}:0<\delta(k,m)\leq\Psi\}.$ (54)
@@ VI Simulations
In this section, we evaluate the performance of the proposed support set recovery methods developed in Section IV and Section V, and compare them with existing methods through numerical simulations. The compared methods and the evaluation measures are presented in Subsection VI-A. We conduct three types of experiments to demonstrate the performance of our sparse recovery methods: 1) on synthetic data over random sbm graphs with known graph filters in Subsection VI-B; 2) for blind identification of graph signals over brain networks in Subsection VI-C; and 3) for recovering the source of the $1854$ Cholera outbreak in London in Subsection VI-D.
VI-A Methods and Performance Measures
@@
  • The Lasso method [23], which is implemented for diffused sparse graph signals in [48];
@@
  • The gmgic method in Algorithm 2;
  • The gbnb method in Algorithm 3 (denoted as G-BNB);
@@
  • The gfoc support set correction method in Algorithm 4, with the estimated support set input from each of the above mentioned methods; The corrected outputs are denoted as 1) omp¯$^{c}$, 2) $\overline{\text{Lasso}}^{c}$, 3) $\overline{\ell_{1}-BnB}^{c}$, 4) gmgic¯$^{c}$, 5) $\overline{\text{G-BNB}}^{c}$, 6) gic¯$^{c}$.
In practice, the Lasso and $\ell_{1}$-bnb methods directly solve the sparse recovery problem in (14). As shown in Section II, this is equivalent to solving (16). The tuning factors for the different methods are stated as follows. The Lasso and $\ell_{1}$-bnb regularization parameter is set to $0.01$. Additionally, the gic penalty function, ${\rho{\lvert{\Omega}\rvert}}$, used for the gmgic, gic, G-BNB, gic, and gfoc methods, is chosen to be the aic penalty function [64], which demonstrated efficacy across a variety of settings we evaluated. Hence, we set ${\rho{\lvert{\Omega}\rvert}}=2{\lvert{\Omega}\rvert}$. For the gmgic method, the pre-screening threshold $\zeta$ in (37) is set to $\sigma_{n}$. For each presented scenario, the performance has been evaluated by at least $10^{3}$ Monte Carlo simulations.
The following performance measures are used in the following to evaluate the different methods:
  • The support set recovery is measured by the F-score classification metric [77]. The F-score compares between the the true support set, ${\Omega}_{t}$, and the estimated support, $\hat{{\Omega}}$, which is obtained by the different methods above. The F-score is given by
@@
    where $t_{p}\mathrel{{}\mathop{=}\limits^{\triangle}}{\lvert{\Omega}_{t}\cap\hat{{\Omega}}\rvert}$ is the number of true-positives, $f_{p}\mathrel{{}\mathop{=}\limits^{\triangle}}{\lvert\{{\mathcal{V}}\setminus{\Omega}_{t}\}\cap\hat{{\Omega}}\rvert}$ is the number of false-positives, and $f_{n}\mathrel{{}\mathop{=}\limits^{\triangle}}{\lvert{\Omega}_{t}\cap\{{\mathcal{V}}\setminus\hat{{\Omega}}\}\rvert}$ is the number of false-negatives. The F-score is between 00 and $1$, where $1$ indicates perfect support set recovery, i.e. $\hat{{\Omega}}={\Omega}$.
  • The sparse signal recovery accuracy is measured using the mse between the true graph signal, ${{\bf{x}}}$, and the estimated sparse graph signal, $\hat{{{\bf{x}}}}$: $MSE({{\bf{x}}},\hat{{{\bf{x}}}})=\left\|{{\bf{x}}}-\hat{{{\bf{x}}}}\right\|_{2}^{2},$ where both ${{\bf{x}}}$ and $\hat{{{\bf{x}}}}$ are normalized. In practice, the sparse signal is recovered by substituting $\hat{{\Omega}}$ in (15).
  • The computational complexity is evaluated by the average run-time of the algorithms, where the simulations have been conducted using Matlab on two Intel(R) Xeon(R) CPU E5-2660 v4 @ 2.00 GHz processors (Subsection VI-B) or using Matlab on the Apple M1 pro chip (Subsections VI-C and VI-D).
Refer to caption(a)
Refer to caption(b)
Refer to caption(c)
Refer to caption(d)
Refer to caption(e)
Refer to caption(f)
Refer to caption(g)
Refer to caption(h)
Refer to caption(i)
Figure 4: SBM graph model: The F-score ((a)-(c)), MSE ((d))-(f)), and run-time ((g)-(i)) of the different methods are compared versus the graph filter degree ($\Psi$), SNR, and graph size.
VI-B Test Case A - Synthetic Data Set
In this subsection, we consider the support set recovery problem in (16), where the sparse dictionary matrix is selected as the graph filter defined by (6). The gso used to construct the graph filter is the graph Laplacian matrix [3], i.e. ${\bf{S}}={\bf{L}}$, which is normalized such that its largest eigenvalue is $12$. The graph filter coefficients are set to $1$, $h_{j}=1$, $j=0,\ldots{\Psi}$. Additionally, the measurement vector ${{\bf{y}}}$ follows the model in (11). The underlying graph is modeled as an sbm graph with ${K}$ clusters, where each cluster contains $N=70$ nodes [78]. The probability of connection between any two nodes within the same cluster is $\frac{6}{N}$. In addition, consecutive clusters are connected by $2$ link nodes from each cluster. Unless stated otherwise, in the following simulations the sbm is composed of $2$ clusters and the support set elements in ${\Omega}$, where ${\lvert{\Omega}\rvert}=4$, are selected from one cluster using one of the following two cases:
@@ VI-B Test Case A - Synthetic Data Set
    Scenario 1 (Mixed Dispersed-Localized Support): In this scenario, two nodes are randomly selected from the cluster, each has one of its neighbors added to the support set. This process is repeated until four different nodes are selected, which results in two groups of four distinct nodes. This scenario describes a situation where the non-zero elements are clustered in two separated groups, where the nodes within each group are in close proximity.